# 3M-Diffusion: Latent Multi-Modal Diffusion for Text-Guided Generation of Molecular Graphs

### v1

@@ 4.2 Multi-modal Molecule Latent Diffusion
We then introduce latent diffusion for conditional molecular graph generation, a method that leverages the aligned molecule-text latent space of encoder-decoder discussed above, enabling us preserve the crucial molecular properties and aligning the textual description of $\mathcal{G}$. Specifically, we learn a better probabilistic mapping from the text to the aligned moelcule latent space, and thereby generate a higher-quality and more diverse molecules with more semantic consistency conforming to the conditional text inputs.
Specifically, the denoising network $\hat{\mathbf{z}}_{\theta}$ focuses on generating latent graph representation $\mathbf{z}$ conditioned on the text representation $\mathbf{c}$ from the text encoder $E_{t}$. The objective is:
$\small\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t,\mathbf{c}\right)-\mathbf{z}\right\|_{2}^{2}\right],$ (6)
@@
In addition, we improve sample quality through classifier-free guidance (Ho & Salimans, 2022), jointly training an unconditional network, $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)$, and a conditional network, $\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)$. Conditioning information is randomly dropped with a probability of $p=0.1$ during training. When conditioning information is dropped, we replace the embedded source text with a learnable embedding.
@@ 4.3 Training and Inference
Training. We now describe the training procedure for 3M-Diffusion. While the learning objectives for graph encoder-decoder and multi-modal latent diffusion have already been specified by Equations (5) and (6), it is unclear whether the two components should be trained sequentially
Previous research on latent diffusion models for image generation indicates that a two-stage training strategy often results in superior performance (Rombach et al., 2022). Hence, we adopt a similar two-stage strategy. Specifically, our first stage is to train the encoder-decoder using Equation (5), and the second stage is to train multi-modal molecule latent diffusion with the fixed encoder-decoder from stage one.
@@
$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\tilde{\mathbf{z}}_{t}$ (8)
$+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$
where the sampling chain is initialized from Gaussian prior $\mathbf{z}_{T}\sim p\left(z_{T}\right)=\mathcal{N}\left(\mathbf{z}_{T};\mathbf{0},\bm{I}\right)$. The detailed training and inference algorithms of 3M-diffusion are in Appendix B.
@@
5 Experiment
5.1 Experimental Settings
Table 3: Statistics of all datasets.
Dataset #Training #Validation #Test
ChEBI-20 15,409 1,971 1,965
PubChem 6,912 571 1,162
PCDes 7,474 1,051 2,136
MoMu 7,474 1,051 4,554
@@ 5.1 Experimental Settings
Datasets. In our experiments, we use PubChem (Liu et al., 2023), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022) and Momu (Su et al., 2022) datasets. Following the evaluation methodology adopted with datasets (ZINC250K (Irwin et al., 2012) and QM9 (Blum & Reymond, 2009; Rupp et al., 2012)) used in previous studies of automated generation of molecular structures, we limit our training, validation, and test data to molecules with fewer than 30 atoms. Statistics of datasets are shown in Table 3 and detailed descriptions are given in the Appendix C.1.
@@
Evaluation. We evaluate the model’s performance on two tasks: text-guided and unconditional molecule generation. Following previous works on text-guided molecule generation, we compare our 3M-Diffusion the SOTA baselines (Edwards et al., 2022; Christofidellis et al., 2023). Text-guided molecule generation measures the model’s capacity to generate chemically valid, structurally diverse, novel molecules that match the textual description provided. Based on previous work (Chen et al., 2021; Fu et al., 2021; Wang et al., 2022), we adopt the following metrics for this task: Similarity, Novelty, Diversity, Validity. We call a predicted molecular graph, specifically one that is chemically valid ($g^{\prime}$), qualified with the ground truth graph $g$ if $f(g,g^{\prime})>0.5$, where $f$ measures the MACCS (Durant et al., 2002) cosine similarity between the two. If the $f(g,g^{\prime})$ similarity between $g^{\prime}$ and $g$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $g$ novel. Specifically, (i) Similarity: the percentage of qualified molecules out of 20K molecules; (ii) Novelty: the percentage of novel molecules out of all qualified molecules; (iii) Diversity: the average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules. (iv) Validity: the percentage of generated molecules that are chemically valid out of all molecules.
@@
We report the results for unconditional molecule generation tasks using the GuacaMol benchmarks (Brown et al., 2019) to assess the model’s capability to learn molecular data distribution and generate realistic, diverse molecules. Specifically, Uniqueness gauges the ratio of unique molecules among the generated ones. Novelty measures the models’ capacity to generate molecules not present in the training set. KL Divergence assesses the proximity between the distributions of various physicochemical properties in the training set and the generated molecules. Fréchet ChemNet Distance (FCD) computes the proximity of two sets of molecules based on their hidden representations in ChemNet (Preuer et al., 2018). Each metric is normalized to a range of 0 to 1, with a higher value indicating better performance.
@@
Baselines. For text guided molecule generation, we compare 3M-Diffusion with the following baselines: MolT5 (Edwards et al., 2022) and ChemT5 (Christofidellis et al., 2023). We consider three representative baselines for unconditional molecule generation based on Variational Autoencoder, including BwR (Diamant et al., 2023), HierVAE (Jin et al., 2020) and PS-VAE (Kong et al., 2022). The descriptions and details of the baselines are given in Appendix C.1.
Setup. We randomly initialize the model parameters and use the Adam optimizer to train the model. HierVAE is used as our variational autoencoder for all of the following experiments. For baselines of unconditional generation that did not report results in these datasets, we reproduce the results using the official code of the authors. To ensure a fair comparison across all methods for unconditional generation, we select the best hyperparameter configuration solely based on the loss on the training set. We evaluate the performance of baselines for text-guided molecule generation using the released pre-trained parameters.
Figure 3: Molecules generated unconditionally by 3M-diffusion from the prior distribution $\mathcal{N}(\mathbf{0},\bm{I})$.
5.2 Text-guided Molecule Generation
Table 1 and 2 report Similarity, Novelty, Diversity, and Validity for text-guided molecule generation on four datasets. All of these metrics are computed as averages, with each metric being calculated based on 5 samples generated from each textual description. Our model demonstrates competitive results in Similarity and notably higher values in the Diversity and Novelty. Specifically, 3M-Diffusion achieves 146.27% (Novelty of PCDes), 130.04% (Diversity of PCDes) relative improvement over the second-best method with maintaining semantics in the textual prompt. This suggests that transformer-based models, pre-trained on a large dataset, excel in predicting molecules based on textual descriptions. However, our model can generate a diverse and novel collection of molecular graphs that match the textual descriptions provided as input. Also, as Figure  4 shows, our model can generate molecules faster than transformer-based methods.
@@ 5.3 Unconditional Molecule Generation
Table 4 shows the results of unconditional molecule generation on PubChem and ChEBI-20. 3M-Diffusion consistently outperforms or achieves competitive results across all four metrics on both datasets, indicating its ability to generate realistic molecules while avoiding overfitting the training data. Furthermore, 3M-Diffusion consistently outperforms HierVAE when both methods use the same decoder. This underscores the clear advantage of 3M-Diffusion in generating a diverse and novel assortment of molecular graphs. As shown in Figure 3, our sampled molecules present rich variety and structural complexity. We have included more sampled molecules from the prior distribution in Appendix D.4.
Figure 4: The average inference time comparison of molecule conditional generation on ChEBI-20 dataset.
5.4 Ablation Study
We conduct ablation studies to assess the impact of different model designs, including joint training of the diffusion model and Variational Autoencoder (denoted as ”Without two stage training”) and training the model without aligning the representation space of graphs and text using $\mathcal{L}_{\text{con}}$ in Equation (4) (denoted as ”Without representation alignment”). Additionally, we perform experiments without using either of these configurations, labeled as ”Without two stage training & Without representation alignment” in Table 5. Initially, training the aligned encoder-decoder alongside Multi-modal Molecule Latent Diffusion jointly yielded comparable or superior performance in terms of diversity and novelty. However, there was a noticeable deficiency in its ability to learn align its representation of molecular graphs with that of their textual descriptions. Training 3M-Diffusion without the aligned encoder-decoder results in substantial drop in performance relative to 3M-Diffusion. This result underscores the importance of aligning the representation of textual descriptions of molecules with that of molecular graphs. Without using both design, Similarity suffers a substantial decline to 10.24, rendering it unsuitable for real-world applications. The full 3M-Diffusion model (last row) achieves the best performance with respect to all of the metrics. This confirms the suitability of the two-step training approach for generating diverse, novel set of molecules based on their textual descriptions.
Table 5: Results of ablation study with different model designs of 3M-diffusion for the Pubchem dataset.
@@ 5.5 Qualitative Analysis and Case Study
Figure 5: Molecules generated conditionally on input text by 3M-Diffusion on the ChEBI-20 dataset.
@@
Property Conditional Molecule Generation. In traditional molecule generation models, the generation process is often conditioned on specific molecular properties such as logP (Kusner et al., 2017) and QED. logP assesses a molecule’s solubility, while QED evaluates its drug likeness (Bickerton et al., 2012). Previous methods have explored ways to optimize their models further in order to generate molecules with desired properties. However, in the case of our model, a different approach is taken. Our model utilizes designed prompts to realize the conditional molecule generation process, as illustrated in Figure 2 for assessing molecule solubility. In particular, we generated 10 samples for each prompt and subsequently select the top 5 samples based on the desired property values. For MolT5-large, we adjust the temperature of the transformer in an attempt to encourage diversity while ensuring that invalid samples are controlled during the sampling process. We applied the same methodology used in our previous experiments to maintain consistency across the comparative analysis, as shown in Figure 2. Compared with the MolT5-large model, we find that 3M-Diffusion excels in generating a diverse range of molecules while achieving improved desired properties, as evidenced by higher logP values Figure 2. This underscores our model’s capacity to produce molecules that are not only novel and diverse but also match the textual description provided. This unique approach distinguishes 3M-Diffusion from prior methods for molecular generation, presenting a fresh and promising perspective on molecular design. For additional experimental details and insights regarding prompts related to solubility and drug likeness, please refer to the Appendix D.1.
@@
Qualitative Comparison. In Figure 2, we introduce more specific prompts related to molecular categories, such as ”The molecule is a member of the class of flavones.” These prompts bias the model, in favor of generating useful drug candidates. Specifically, we choose a reference molecule from the training dataset that matches the given textual prompt. Subsequently, we generate ten additional molecules and select the top five molecules that exhibit the highest structural similarity (calculated by Rascal algorithm (Raymond et al., 2002)) to the reference molecule. These selected molecules are shown in Figure 2.Comparisong of our model with MolT5-large shows that our model generates more diverse molecules that exhibit a higher degree of structural similarity to the reference molecule. Simultaneously, it also produces novel molecular graphs. We also include more prompt examples in Appendix D.2.
@@
Case Study. In Figure 5, we choose a set of examples from the ChEBI-20 dataset to visually illustrate the structural similarity between the molecules generated by 3M-Diffusion and the ground truth molecules corresponding to the given text descriptions. Specifically, we show the largest substructure that appears in generated and ground truth molecules by RDKit. 3M-Diffusion demonstrates the capability to generate molecules that share a common substructure, as illustrated in Figure 5. Simultaneously, it has the flexibility to explore novel components or structures, thereby achieving a balance between generating similar structures and introducing novel molecules. This observation further confirms the ability of our model in generating a diverse collection of novel molecular structures that match the textual description provided. We show more results in Appendix D.3.

### v2

@@ 4.2 Multi-modal Molecule Latent Diffusion
We next introduce latent diffusion for conditional molecular graph generation, a method that leverages the aligned molecule-text latent space of encoder-decoder described above, enabling us to preserve the crucial molecular properties and aligning the textual description of $\mathcal{G}$. Specifically, we learn a probabilistic mapping from the text to the text aligned molecular structure latent space, and thus generate molecular structures that match the conditional text inputs.
The denoising network $\hat{\mathbf{z}}_{\theta}$ focuses on generating latent graph representation $\mathbf{z}$ conditioned on the text representation $\mathbf{c}$ from the LLM encoder $E_{t}$. The objective to be optimized is:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\lambda_{t}\left\|\hat{\mathbf{z}}_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}},t,\mathbf{c}\right)-\mathbf{z}\right\|_{2}^{2}\right],$ (6)
@@
Additionally, to improve sample quality we take advantage of classifier-free guidance (Ho & Salimans, 2022), to jointly train an unconditional network, $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)$, and a conditional network, $\hat{\mathbf{z}}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{c}\right)$. Conditioning information is randomly dropped with a probability of $p=0.1$ during training. When conditioning information is dropped, we replace the embedded source text with its embedding.
@@ 4.3 Training and Inference
Training. We proceed to describe the training procedure for 3M-Diffusion. While the learning objectives for graph encoder-decoder and multi-modal latent diffusion have already been specified by Equations (5) and (6), it is unclear whether the two components should be trained sequentially. Previous research on latent diffusion models for image generation indicates that a two-stage training strategy often results in superior performance (Rombach et al., 2022). Hence, we adopt a similar two-stage strategy. The first stage consists of training an encoder-decoder pair using Equation (5), and the second stage consists of training a multi-modal molecule latent diffusion with the fixed encoder-decoder from the first stage.
@@
$\mathbf{z}_{t-1}$ $=\tfrac{\sqrt{\alpha_{t-1}}\left(1-\alpha_{t\mid t-1}\right)}{1-\alpha_{t}}\tilde{\mathbf{z}}_{t}+\tfrac{\sqrt{\alpha_{t\mid t-1}}\left(1-\alpha_{t-1}\right)}{1-\alpha_{t}}\mathbf{z}_{t}+\sigma(t-1,t){\bm{\epsilon}},$ (8)
where the sampling chain is initialized from Gaussian prior $\mathbf{z}_{T}\sim p\left(z_{T}\right)=\mathcal{N}\left(\mathbf{z}_{T};\mathbf{0},\bm{I}\right)$. The detailed training and inference algorithms of 3M-diffusion are provided in Appendix B.
Table 1: Quantitative comparison of conditional generation on the PCDes and MoMu. 3M-Diffusion outperforms other SOTA methods in terms of novelty, diversity, and validity metrics by a large margin, while maintaining a good similarity metric.
@@
5 Experiments and Results
5.1 Experimental Setup
@@
Datasets. We use PubChem (Liu et al., 2023), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022) and Momu (Su et al., 2022) datasets. Following the evaluation methodology adopted with datasets (ZINC250K (Irwin et al., 2012) and QM9 (Blum & Reymond, 2009; Rupp et al., 2012)) used in previous studies of molecular structure generation, we limit our training, validation, and test data to molecules with fewer than 30 atoms. Statistics of datasets are shown in Table 5 and detailed descriptions are given in the Appendix C.1.
@@
Evaluation. We evaluate the model’s performance on two tasks: text-guided and unconditional molecule generation. Following previous works on text-guided molecule generation, we compare our 3M-Diffusion the SOTA baselines (Edwards et al., 2022; Christofidellis et al., 2023; Fang et al., 2023a). Text-guided molecule generation measures the model’s capacity to generate chemically valid, structurally diverse, novel molecules that match the textual description provided. As in previous work (Chen et al., 2021; Fu et al., 2021; Wang et al., 2022), we adopt the following metrics for this task: Similarity, Novelty, Diversity, Validity. We call a predicted molecular structure, specifically one that is chemically valid ($g^{\prime}$), qualified with the ground truth graph $g$ if $f(g,g^{\prime})>0.5$, where $f$ measures the MACCS (Durant et al., 2002) cosine similarity between the two. If the $f(g,g^{\prime})$ similarity between $g^{\prime}$ and $g$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $g$ novel. Specifically, (i) Similarity: the percentage of qualified molecules out of 20K molecules; (ii) Novelty: the percentage of novel molecules out of all qualified molecules; (iii) Diversity: the average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules. (iv) Validity: the percentage of generated molecules that are chemically valid out of all molecules.
@@
We report the results for unconditional molecule generation tasks using the GuacaMol benchmarks (Brown et al., 2019) to assess the model’s ability to learn a generative model of molecular structures that can be used to generate realistic, diverse molecules. Specifically, Uniqueness gauges the ratio of unique molecules among the generated ones. Novelty measures the models’ capacity to generate molecules not present in the training set. KL Divergence assesses the proximity between the distributions of various physicochemical properties in the training set and the generated molecules. Fréchet ChemNet Distance (FCD) computes the proximity of two sets of molecules based on their hidden representations in ChemNet (Preuer et al., 2018). Each metric is normalized to a range of 0 to 1, with a higher value indicating better performance.
Baselines. For text guided molecule generation, we compare 3M-Diffusion with the following baselines: MolT5 (Edwards et al., 2022), ChemT5 (Christofidellis et al., 2023) and Mol-Instruction (Fang et al., 2023a). We consider seven representative baselines for unconditional molecule generation based on Variational Autoencoder, including CharRNN (Segler et al., 2018), VAE (Kingma & Welling, 2013), AAE (Makhzani et al., 2015), LatentGAN (Prykhodko et al., 2019) BwR (Diamant et al., 2023), HierVAE (Jin et al., 2020) and PS-VAE (Kong et al., 2022). The details of the baselines are given in Appendix C.1.
Setup. We randomly initialize the model parameters and use the Adam optimizer to train the model. HierVAE is used as our variational autoencoder for all of the following experiments. For baselines of unconditional generation that did not report results in these datasets, we reproduce the results using the official code made available by the authors. To ensure a fair comparison across all methods for unconditional generation, we select the best hyperparameter configuration solely based on the loss on the training set. We evaluate the performance of baselines for text-guided molecule generation using the pretrained parameters made available by the authors of the respective studies.
5.2 Text-guided Molecule Generation
Figure 3: Inference time comparison for conditional molecule generation on ChEBI-20.
Table 1 and 2 report Similarity, Novelty, Diversity, and Validity for text-guided molecule generation on four datasets. Each metric is computed as an average over 5 structures generated from a given textual description. We see that 3M-Diffusion is competitive with SOTA baselines with respect to Similarity and outperforms them substantially with respect to Diversity and Novelty. Specifically, 3M-Diffusion achieves 146.27% (Novelty of PCDes), 130.04% (Diversity of PCDes) relative improvement over the second-best SOTA baseline while matching the textual prompt. We conclude that while the transformer based models that are pretrained on a large dataset excel in producing molecular structures that match the textual descriptions provided, 3M-Diffusion outperforms them with respect to the diversity and novelty of the molecular structures without sacrificing the match between the textual descriptions provided as input and the molecular structures generated as output. Also, as Figure  3 shows, 3M-Diffusion enjoys a substantial advantage in terms of its speed over transformer-based methods (See Appendix D.1 for details).
@@ 5.3 Unconditional Molecule Generation
Table 3 shows the results of unconditional molecule generation on PubChem and ChEBI-20. 3M-Diffusion consistently outperforms or matches the performance of SOTA baselines across all four metrics on both datasets, suggesting that it is able to generate realistic, diverse, and novel molecules without overfitting the training data. Furthermore, 3M-Diffusion consistently outperforms HierVAE when both methods use the same decoder. This underscores the clear advantage of 3M-Diffusion in generating a diverse and novel molecular graphs (See Figure 10 for 50 samples from the learned distribution).
5.4 Ablation Studies
We conduct ablation studies to assess the impact of different model designs, by joint training of the diffusion model and Variational Autoencoder (”without two-stage training”) and training the model without aligning the representation space of graphs and text using $\mathcal{L}_{\text{con}}$ in Equation (4) ( ”without representation alignment”). Additionally, we perform experiments without both two-stage training and representation alignment (”without two-stage training & representation alignment”) in Table 4. We find that joint training the aligned encoder-decoder alongside Multi-modal Molecule Latent Diffusion achieves comparable or superior performance relative to 3M-Diffusion in terms of diversity and novelty, but at the expense of reduced quality of match between the molecular structures produced and the textual descriptions provided. We find that training the model without the aligned encoder-decoder results in substantial drop in performance relative to 3M-Diffusion, underscoring the importance of aligning the representation of textual descriptions of molecules with that of molecular structures. Ablation of both two-stage training and representation alignment components leads to a substantial drop in performance (Similarity of 10.24), rendering the resulting model unsuitable for real-world applications. The full 3M-Diffusion model (last row) achieves the best performance with respect to all of the metrics. This confirms the importance of two-step training and representation alignment in training a model that can generate diverse, novel set of molecular structures from their textual descriptions.
Table 4: Results of ablation study with different model designs of 3M-diffusion for Pubchem.
@@ 5.5 Qualitative Analysis and Case Study
Property Conditional Molecule Generation. In traditional molecule generation models, the generation process is often conditioned on specific molecular properties such as logP (Kusner et al., 2017) and QED. logP assesses a molecule’s solubility, while QED evaluates its drug likeness (Bickerton et al., 2012). Previous methods have explored ways to optimize such models further in order to generate molecules with desired properties. In contrast, 3M-Diffusion utilizes textual prompts for conditional molecule generation (See Figure 2). In particular, we generated 10 samples for each prompt and subsequently select the top 5 samples based on the value of the desired property. For MolT5-large, we adjust the temperature of the transformer in an attempt to encourage diversity while ensuring that invalid samples are controlled during the sampling process. We applied the same methodology used in our previous experiments to maintain consistency across the comparative analysis, as shown in Figure 2. When compared to the MolT5-large model, we find that 3M-Diffusion excels in generating a diverse range of molecules that also score better with respect to the desired properties, as evidenced by higher logP values Figure 2. This underscores our 3M-Diffusion’s ability to produce molecules that are not only novel and diverse but also match the textual description provided. This unique approach distinguishes 3M-Diffusion from prior methods for molecular generation, presenting a fresh and promising perspective on molecular design. Additional experimental details and insights regarding prompts related to solubility and drug likeness are provided in Appendix D.2.
Figure 4: Molecules generated conditionally on input text by 3M-Diffusion on ChEBI-20.
@@
Qualitative Comparison. In Figure 2, we introduce more specific prompts related to molecular categories, such as ”The molecule is a member of the class of flavones.” These prompts bias the model in favor of generating useful drug candidates. Specifically, we choose a reference molecule from the training dataset that matches the given textual prompt. Subsequently, we generate ten additional molecules and select the top five molecules that exhibit the highest structural similarity (calculated by Rascal algorithm (Raymond et al., 2002)) to the reference molecule. These selected molecules are shown in Figure 2.Comparison of 3M-Diffusion with MolT5-large shows that 3M-Diffusion generates more diverse molecules that exhibit a higher degree of structural similarity with the reference molecule while producing novel molecular structures. Additional examples of prompts are provided in Appendix D.3.
@@
Case Study. In Figure 4, we choose a set of examples from the ChEBI-20 dataset to visually illustrate the structural similarity between the molecules generated by 3M-Diffusion and the ground truth structures corresponding to the given textual descriptions. Specifically, we show the largest substructure shared by the generated and ground truth molecules (using RDkit (Landrum, 2010)). 3M-Diffusion demonstrates the capability to generate molecules that include the relevant substructure, as illustrated in Figure 4. Simultaneously, it displays the flexibility to generate molecules with novel substructures, thereby striking a balance between generating structures that are similar to those represented in the training data and suggesting novel structures that match the textual description provided. We include additional results in Appendix D.4.